OpenStudy (anonymous):

Last question on my test need it right for I can pass help please. Determine the equation of the line in point-slope form, that runs through points (2,7) and (-2,1) A: y-7=3/2 (x-2) B: y+7=3/2 (x+2) C: y-7=3/2 x-2

OpenStudy (kamibug):

A

OpenStudy (anonymous):

Im sorry to ask but u sure? cause im sorry im really realy trying to get this right if I miss I fail my exam. Sorry for questioning u.

OpenStudy (kamibug):

Lol, it's okay, don't worry. I am sure. Point-Intercept form is y-y1=m(x-x1) First you need to find the slope using the slope formula: m = y2-y1/x2-x1 1-7/-2-2 = 6/4 = 3/2 Then just plug in either of the points. Here they want you to use (2, 7) y-7=3/2(x-2)

OpenStudy (anonymous):

Thank you soo much thank you :)

OpenStudy (kamibug):

Sure :)